# Unlearning Reveals the Influential Training Data of Language Models

### v1

@@ Results
Table 2 presents the influence of training datasets estimated by each method and the ground-truth influence measured by leave-dataset-out. The ground-truth indicates that datasets 1 and 2 (tasks similar to the test task) are more influential than datasets 3 and 4 (tasks dissimilar to the test task). The influence estimated by UnTrac and UnTrac-Inv aligns well with the ground-truth influence. All other methods assess the influence of dataset 4 as lower than that of other datasets. However, they tend to overestimate the influence of dataset 3 or underestimate it for datasets 1 and 2.
@@ 5.2 Tracing Influential Pretraining Corpora
Regarding test datasets, we use three datasets: Toxigen Hartvigsen et al. (2022), WinoBias Zhao et al. (2018), and TruthfulQA Lin et al. (2022). ToxiGen collects machine-generated toxic language against 13 minority groups. Winobias contains sentences with entities corresponding to people referred to by their occupation and their pronoun genders. TruthfulQA comprises questions across several categories and their corresponding untruthful answers. We measure the negative log-likelihood of toxic language, biased text, and false answers to compute the influences.
@@
Table 3: Pearson correlation coefficient between the influence estimated by each method and the ground-truth influence computed by leave-dataset-out. Each figure denotes the mean and standard deviation across four runs. For each run, we use different examples randomly sampled from the training dataset to compute its influence.
Figure 3: Pearson correlation coefficient between the ground-truth influence and the influence estimated by UnTrac (left) and UnTrac-Inv (right) over unlearning epochs. The line denotes the average across four runs, and the shaded area corresponds to 95% confidence region.
@@ Results
Table 3 shows the Pearson correlation coefficient between the estimated influence and the ground-truth assessed by leave-dataset-out. [^6] The same tendency is confirmed when Spearman’s rank correlation coefficient is used as a metric (Appendix A.3).
Across all datasets and settings, the estimated influence by UnTrac and UnTrac-Inv correlates well with the ground-truth. GradCos, GradDot, and HIF (Arnoldi) perform well on Winobias. However, they show lower performance on Toxigen and TruthfulQA specifically when the dataset sizes are unbalanced. The performance of HIF (LISSA) is unstable, as indicated by the high variance in its score. While TracIn achieves a relatively higher correlation with equally sized training datasets, its performance declines when the sizes are different.
These results indicate that our methods can trace the influence of pretraining corpora throughout the whole training process. Our methods maintain robustness under varying dataset proportions.
@@ 6 Discussion
In this section, we explore how hyperparameters affect the performance of our methods. Following the experiment in Section 5.2, we compute the influence of the pretraining dataset on Toxigen under various hyperparameter settings. By monitoring the Pearson correlation coefficient between the estimated influence and ground-truth, we investigate the hyperparameter sensitivity of our methods.
@@ 6.1 Sensitivity to Epoch & Batch Size
Optimizer
SGD
SGD w/momentum
RMSProp
Adam
Adafactor
UnTrac
-0.147±0.014
-0.239±0.011
0.418±0.063
0.419±0.063
0.345±0.179
UnTrac-Inv
-0.100±0.069
-0.099±0.070
-0.231±0.012
0.376±0.008
0.313±0.003
Learning Rate
5e-06
1e-05
5e-05
1e-04
5e-04
UnTrac
-0.127±0.302
0.312±0.311
0.419±0.063
0.377±0.040
0.329±0.015
UnTrac-Inv
0.100±0.084
0.197±0.067
0.376±0.008
0.137±0.019
0.027±0.015
@@ 6.2 Sensitivity to Optimizer & Learning Rate
Table 4 (top) shows the performance with various optimizers. We use SGD, SGD with momentum ($\mathrm{momentum}\!=\!\mathrm{dampening}\!=\!0.9$), RMSProp (Tieleman et al., 2012; $\alpha=0.99$), Adam, and Adafactor Shazeer and Stern (2018). UnTrac performs well when RMSProp, Adam, and Adafactor are used, indicating that a preconditioner plays an important role in unlearning. The same trend can be seen for UnTrac-Inv, though it performs worse when RMSProp is used.
@@
Table 4 (bottom) presents the performance with several learning rates. UnTrac performs well across various learning rates when higher learning rates are used. With lower learning rates, UnTrac does not converge and performs unstably. In contrast, UnTrac-Inv is somewhat sensitive to the choice of learning rate; however, the estimated influence correlates positively with ground-truth influence on any learning rate.
@@ 7 Related Work
Which training data affects the model’s prediction? The importance of this question has been rising along with the emergence of LLMs trained on massive text corpora. Hessian-based influence functions Hampel (1974); Koh and Liang (2017) are pioneering work to address this question and are widely used in the context of natural language processing Han et al. (2020); Guo et al. (2021); Yang et al. (2023). As computing inverse Hessian is quite expensive Guo et al. (2021); Schioppa et al. (2022); Grosse et al. (2023), TracIn is another promising approach for training data attribution Pruthi et al. (2020). However, several studies empirically reported that these methods are unstable for deep neural networks Basu et al. (2021); Søgaard (2021).
@@
Unlearning has been studied to remove sensitive data from a trained model Cao and Yang (2015); Ginart et al. (2019); Golatkar et al. (2020); Sekhari et al. (2021); Gupta et al. (2021); Mehta et al. (2022) and has recently been used for LLMs Jang et al. (2023); Wang et al. (2023b).
@@ 8 Conclusion
This paper explored an unlearning approach to trace the influence of training datasets. Our approach, UnTrac, simply unlearns each training dataset and evaluates how the model’s performance changes on a test dataset after unlearning. UnTrac-Inv is a more scalable approach, which unlearns a test dataset and evaluates the unlearned model’s performance on training datasets. Experimental results showed that our methods can trace the influence of training datasets with significantly higher accuracy than existing methods. UnTrac works robustly if we use preconditioned optimizers with high learning rates and a sufficient number of training steps. In contrast, UnTrac-Inv is relatively sensitive to the choice of hyperparameters, such as batch size, learning rate, and the number of training iterations.
As our methods require the same memory footprint as standard training, they can be applied to LLMs as long as we have enough memory space for training. In future work, we hope that the effectiveness of our methods will be investigated with even broader sets of tasks, and that our methods will be helpful in revealing the source of LLM’s emergent abilities, such as chain-of-thought reasoning.
@@ A.1 How to Compute Leave-Dataset-Out
A.2 Relation to Other Influence Functions
The increase in test loss after a single step of unlearning can be approximated as Equation (9).
$\approx\sum_{j=1}^{N^{\prime}}\nabla_{\bm{\theta}}L(z^{\prime}_{j},\bm{\theta}_{0})(\bm{\theta}_{1}-\bm{\theta}_{0})$
Suppose a single training batch contains all training examples. When we use SGD, Equation (9) corresponds to GradDot by substituting $\bm{\theta}_{1}-\bm{\theta}_{0}=\eta_{0}\sum_{i=1}^{N}\nabla_{\bm{\theta}}L(z_{i},\bm{\theta}_{0})$. Similarly, It corresponds to GradCos if we use RMSProp or Adam: $\bm{\theta}_{1}-\bm{\theta}_{0}=\eta_{0}\sum_{i=1}^{N}\nabla_{\bm{\theta}}L(z_{i},\bm{\theta}_{0})/\|\sum_{i=1}^{N}\nabla_{\bm{\theta}}L(z_{i},\bm{\theta}_{0})\|$, and corresponds to HIF if we use Newton’s method: $\bm{\theta}_{1}-\bm{\theta}_{0}=\sum_{i=1}^{N}\bm{H}_{\bm{\theta}}^{-1}\nabla_{\bm{\theta}}L(z_{i},\bm{\theta}_{0})$. In the same way, these influence functions can also be regarded as a special case of UnTrac-Inv.
@@ A.3 Evaluation by Spearman Correlation
Table 5 shows the Spearman’s rank correlation between the estimated influence and the ground-truth influence computed by leave-dataset-out in Section 5.2. The tendency is similar to the result shown in Table 3, which uses the Pearson correlation as an evaluation metric. Across all datasets and settings, the estimated influence by UnTrac and UnTrac-Inv correlates well with the ground-truth influence.

### v2

@@ Results
Table 2 presents the influence of training datasets estimated by each method and the ground-truth influence measured by leave-dataset-out. The ground truth indicates that datasets 1 and 2 (tasks similar to the test task) are more influential than datasets 3 and 4 (tasks irrelevant to the test task). The influence estimated by UnTrac and UnTrac-Inv aligns well with the ground-truth influence. All other methods assess the influence of dataset 4 as lower than that of other datasets. However, they tend to overestimate the influence of dataset 3 or underestimate it for datasets 1 and 2.
@@ 5.2 Tracing Influential Pretraining Corpora
Regarding test datasets, we use three datasets: Toxigen Hartvigsen et al. (2022), WinoBias Zhao et al. (2018), and TruthfulQA Lin et al. (2022). ToxiGen collects machine-generated toxic language against several minority groups. Winobias contains sentences with entities corresponding to people referred to by their occupation and their pronoun genders. TruthfulQA comprises questions across several categories and their corresponding untruthful answers. We measure the negative log-probability of toxic language, biased text, and false answers to compute the influences.
@@
Table 3: Pearson correlation coefficient between the influence estimated by each method and the ground-truth influence computed by leave-dataset-out. Each figure denotes the mean and standard deviation across four runs. For each run, we use different 10,000 examples randomly sampled from the training dataset to compute its influence. The highest score and the scores that lie within its standard deviation are highlighted in bold.
Figure 3: Pearson correlation coefficient between the ground truth and the influence estimated by UnTrac (left) and UnTrac-Inv (right) over unlearning epochs. ToxiGen is used as an evaluation dataset and the size of each training dataset is set to be equal. The line denotes the average across four runs, and the shaded area corresponds to 95% confidence region.
@@ Results
Table 3 shows the Pearson correlation coefficient between the estimated influence and the ground truth assessed by leave-dataset-out. [^5] The same tendency is confirmed when Spearman’s rank correlation coefficient is used as a metric (Appendix A.3). As there are only eight pretraining datasets, the correlation between ground truth and estimated influences cannot be statistically significant. Therefore, we divided the test dataset into multiple subsets and aggregated the correlation for each subset to enhance the reliability of the results. We reported the average across 13 subsets for ToxiGen, 4 subsets for WinoBias, and 9 subsets for TruthfulQA.
Across all datasets and settings, the estimated influence by UnTrac and UnTrac-Inv correlates well with the ground truth. GradCos, GradDot, and HIF (Arnoldi) perform well on Winobias. However, they show lower performance on Toxigen and TruthfulQA specifically when the dataset sizes are unbalanced. The performance of HIF (LISSA) is unstable, as indicated by the high variance in its score. While TracIn achieves a relatively higher correlation with equally sized training datasets, its performance declines when the sizes are different.
We found that the influences estimated by existing methods are not significantly different across training datasets. For WinoBias and TruthfulQA, the ground-truth influence of each pre-training dataset does not differ that much. Thus, the existing methods achieve high correlations on those datasets as well as ours. In contrast, on ToxiGen, the ground-truth influences of pre-training datasets vary significantly; Pile-CC and BookCorpus are more influential than others. Our methods can detect these influential datasets, while other approaches fail, leading to higher performance of our methods on ToxiGen. These results indicate that our methods can robustly trace the influence of pretraining datasets regardless of their proportions.
@@ 6 Discussion
In this section, we explore how hyperparameters affect the performance of our methods. Following the experiment in Section 5.2, we compute the influence of the pretraining dataset on Toxigen under various hyperparameter settings, where the size of each training dataset is set to be equal. By monitoring the Pearson correlation coefficient between the estimated influence and ground truth, we investigate the hyperparameter sensitivity of our methods.
@@ 6.1 Sensitivity to Epoch & Batch Size
Optimizer SGD SGD w/momentum RMSProp Adam Adafactor
UnTrac -0.147±0.014 -0.239±0.011 0.418±0.063 0.419±0.063 0.345±0.179
UnTrac-Inv -0.100±0.069 -0.099±0.070 -0.231±0.012 0.376±0.008 0.313±0.003
Learning Rate 5e-06 1e-05 5e-05 1e-04 5e-04
UnTrac -0.127±0.302 0.312±0.311 0.419±0.063 0.377±0.040 0.329±0.015
UnTrac-Inv 0.100±0.084 0.197±0.067 0.376±0.008 0.137±0.019 0.027±0.015
@@ 6.2 Sensitivity to Optimizer & Learning Rate
Table 4 (top) shows the performance with various optimizers. We use SGD, SGD with momentum ($\mathrm{momentum}\!=\!\mathrm{dampening}\!=\!0.9$), RMSProp (Tieleman et al., 2012; $\alpha=0.99$), Adam, and Adafactor Shazeer and Stern (2018). UnTrac performs well when RMSProp, Adam, and Adafactor are used, indicating that a preconditioner plays an important role in unlearning.
A similar trend can be seen for UnTrac-Inv, though it performs worse when RMSProp is used. We suspect this is because both Adam and Adafactor use a decaying average of gradients, rather than just the gradient for the current batch as RMSProp and SGD do. This makes training less stochastic and has a somewhat similar effect to using larger batches, which is beneficial for UnTrac-Inv.
@@
Table 4 (bottom) presents the performance with several learning rates. UnTrac performs well across various learning rates when using higher learning rates. With lower learning rates, UnTrac does not converge and performs unstably.
In contrast, UnTrac-Inv is somewhat sensitive to the choice of learning rate. Following the discussion in Section 3.2, UnTrac-Inv diverges from UnTrac as unlearning proceeds. As a higher learning rate boosts the divergence, increasing the learning rate does not necessarily lead to higher performance. However, the estimated influence correlates positively with ground truth on any learning rate.
@@ 7 Related Work
Which training data affects the model’s prediction? The importance of this question has been rising along with the emergence of LLMs trained on massive text corpora Akyurek et al. (2022); Longpre et al. (2023); Feng et al. (2023); Han et al. (2023); Grosse et al. (2023). Hessian-based influence functions Hampel (1974); Koh and Liang (2017) are pioneering work to address this question and are widely used in the context of natural language processing Han et al. (2020); Guo et al. (2021); Yang et al. (2023). As computing inverse Hessian is quite expensive Guo et al. (2021); Schioppa et al. (2022); Grosse et al. (2023), TracIn is another promising approach for training data attribution Pruthi et al. (2020). However, several studies empirically reported that these methods are unstable for deep neural networks Basu et al. (2021); Søgaard (2021).
@@
Sampling-based attribution methods, such as Shapley values Ghorbani and Zou (2019), empirical influences Feldman and Zhang (2020), and datamodels Ilyas et al. (2022) are also promising approaches for tracing influential training datasets. However, they commonly require thousands of pairs of training subsets and models trained on the subsets to estimate how the combination of training examples affects the model’s performance. Although TRAK Park et al. (2023), a recent sampling-based approach, requires significantly fewer computational resources than others, it still requires tens of pairs of training subsets and their corresponding trained models, which can be costly as well as leave-dataset-out. Indeed, TRAK was evaluated only on fine-tuning setups in its original paper. On the other hand, we study approaches that are scalable to large pretraining datasets.
Unlearning has been studied to remove undesired data from a trained model Cao and Yang (2015); Golatkar et al. (2020); Sekhari et al. (2021); Gupta et al. (2021) and has recently been used for LLMs Jang et al. (2023); Wang et al. (2023); Yao et al. (2023); Maini et al. (2024); Zhang et al. (2024); Jia et al. (2024).
@@ 8 Conclusion
This paper explored unlearning approaches to trace the influence of training datasets. Our approach, UnTrac, simply unlearns each training dataset and evaluates how the model’s performance changes on a test dataset after unlearning. UnTrac-Inv is a more scalable approach, which unlearns a test dataset and evaluates the unlearned model’s performance on training datasets. Experimental results showed that our methods can trace the influence of training datasets with significantly higher accuracy than existing methods. UnTrac works robustly if we use preconditioned optimizers with high learning rates and a sufficient number of training steps. In contrast, UnTrac-Inv is relatively sensitive to the choice of hyperparameters, such as batch size, learning rate, and the number of training iterations.
As our methods require the same memory footprint as standard training, they can be applied to LLMs as long as we have enough memory space for training. In future work, we hope that the effectiveness of our methods will be investigated with even more large models, and that our methods will be helpful in revealing the source of LLM’s emergent abilities, such as chain-of-thought reasoning.
Limitations
Although our methods are assessed across different model architectures, model sizes, and training setups, further empirical investigations are expected to verify their effectiveness. We do not argue that our methods are immediately applicable in practical use. For real-world applications, we need to validate our methods in various settings and tasks against leave-dataset-out, though leave-dataset-out requires considerable computational cost.
Acknowledgements
MI is partially supported by JST CREST JPMJCR21D1, NEDO JPNP20006, and JSPS KAKENHI 23K16940, Japan. IT is supported by the Dutch National Science Foundation (NWO Vici VI.C.212.053).
@@ A.1 How to Compute Leave-Dataset-Out
A.2 Implementation Details
Our code is implemented with Python v3.8.13, PyTorch v1.12.0 Paszke et al. (2019), and Transformers v4.18.0 Wolf et al. (2020). In Section 5.1, we used the synthetic datasets which are originally created in this study. The numbers used in the output text range from 1 to 256, while the characters and their part of speech are collected from WordNet (Miller, 1994, WordNet 3.0 license). Our study was conducted under the licenses and terms of the scientific artifacts.
Our experiments were conducted with a single NVIDIA A100 (80GB) for each run in Section 5.1 and a single Tesla V100 (16GB) for each run in Section 5.2. Unlearning each training dataset (UnTrac) takes approximately two minutes in Section 5.1 and four hours in Section 5.2. Unlearning each test dataset (UnTrac-Inv) takes approximately one minute in Section 5.1 and 30 minutes in Section 5.2, not including the time required for evaluation.
@@ A.3 Evaluation by Spearman Correlation
Table 5 shows the Spearman’s rank correlation between the estimated influence and the ground-truth influence computed by leave-dataset-out in Section 5.2. The tendency is similar to the result shown in Table 3, which uses the Pearson correlation as an evaluation metric. Across all datasets and settings, the estimated influence by UnTrac and UnTrac-Inv correlates well with the ground truth.